# H-1117 Budapest, Pázmány Péter sétány 1/A, Hungary

### v1

@@ Abstract
Context. The migration of low‑mass planets ($M_{\mathrm{p}}\leq 10\penalty 10000\ M_{\oplus}$) is tightly controlled by the torques exerted by both gas and solids in their natal disks. While canonical models assume a solar solid‑to‑gas mass ratio ($\epsilon\simeq 0.01$) and neglect the back-reaction of solid component of the disk, recent work suggests that enhanced metallicity can radically alter these torques.
@@
Methods. We performed global, two‑dimensional hydrodynamic simulations that treat solid material as a pressureless fluid fully coupled to the gas through drag and include the reciprocal back‑reaction force. The low-mass planet was maintained on a fixed circular orbit, thus we computed static torques. The Stokes number was varied from 0.01 to 10, three surface‑density slopes ($p=0.5,\penalty 10000\ 1.0,\mathrm{and}\penalty 10000\ 1.5$) and three accretion efficiencies ($\eta=0\%,\penalty 10000\ 10\%,\mathrm{and}\penalty 10000\ 100\%$) were explored. Predicted torques, obtained by rescaling canonical $\epsilon=0.01$ results, were compared with direct simulations.
Results. Solid torques scale nearly linearly with $\epsilon$, but gas torques deviate by $50-100\%$ and can even reverse sign for $\mathrm{St}\leq 1$ in $\epsilon=0.1$ disks. These discrepancies arise from strong, feedback‑driven, asymmetric gas perturbations in the co‑orbital region, amplified by rapid planetary accretion. Accurate total torques are recovered only for $\mathrm{St}\geq 3$, independent of $\epsilon$ or $\eta$; for $\mathrm{St}\leq 2$ the linear prescription systematically overestimates the magnitude, sometimes predicting the wrong sign.
@@ 1 Introduction
It is well established that planets undergo changes in their semi-major axes due to gravitational interactions with their natal protoplanetary disk. In particular, low-mass planets (with masses $\lesssim 10\penalty 10000\ M_{\oplus}$) are generally subject to inward migration in locally isothermal disks as a result of disk torques (Tanaka et al., 2002). Consequently, planets that remain embedded in the disk during its lifetime are at risk of spiraling into their host star.
@@
To counteract this potentially catastrophic inward migration, several mechanisms have been proposed, including the concept of planetary migration traps (Lyra et al., 2008; Sándor et al., 2011; Guilera and Sándor, 2017; Pierens and Raymond, 2024) and localized temperature perturbations within the disk (Lyra et al., 2010). For comprehensive reviews of the physical processes governing planetary migration (see, e.g., Kley and Nelson, 2012; Raymond et al., 2020; Paardekooper et al., 2023.
@@
In recent years, numerical simulations of planet-disk interactions have introduced a new paradigm, emphasizing the pivotal role of the disk’s solid component in altering – and in some cases reversing – the direction of planetary migration. It has been demonstrated that a pronounced asymmetry in the spatial distribution of solids near the planet can generate an additional torque, which may reach magnitudes comparable to that of the gas torque (Benítez-Llambay and Pessah, 2018). Moreover, planetary accretion can further enhance this solid-induced torque, thereby amplifying its impact on the planet’s orbital evolution (Regály, 2020; Chrenko et al., 2024; Pierens, 2023). When these findings are applied within a global disk model that incorporates viscous accretion and X-ray photoevaporation of the gaseous component, along with dust evolution processes, it becomes evident that the torque exerted by the solid phase can significantly influence migration behavior (Guilera et al., 2023). In particular, such models reveal that planets forming near the water-ice line may undergo substantial outward migration driven by solid torques (Guilera et al., 2025).
@@
The back-reaction of solid material on the gas is often neglected in numerical simulations of planet-disk interactions. Nonetheless, a growing body of work has underscored the significance of this feedback mechanism (Gonzalez et al., 2017; Dipierro et al., 2018; Gonzalez et al., 2018; Hsieh and Lin, 2020). Our recent study has shown that back-reaction from the solid component can substantially perturb the gas distribution in the vicinity of the planet, to the extent that the gas torque becomes positive, even under locally isothermal conditions (Regály et al., 2025). In addition, a recent three-dimensional hydrodynamic simulations demonstrated that this solid back-reaction can excite buoyancy resonances in a globally isothermal disk that would otherwise lack such instabilities (Chametla et al., 2025b).
@@
In the canonical scenario, the solid-to-gas mass ratio in protoplanetary disks is approximately one percent (e.g., Williams and Cieza, 2011). It is commonly assumed that the torque exerted by the solid component scales linearly with the total solid mass in the disk (see, e.g., Benítez-Llambay and Pessah (2018)). This assumption may break down when the back-reaction from the solid component becomes sufficiently strong to perturb the gas dynamics. Such perturbations can significantly affect the evolution of the solid component itself, resulting in a nonlinear and coupled torque response.
@@
Hsieh and Lin (2020) showed that higher-metallicity disks can produce solid torques strong enough to alter the migration direction of low-mass planets or to induce stochastic migration behavior. The stochastic nature of migration is associated with the excitation of small-scale vortices produced in inviscid simulations. Their work also accounted for the dynamical back-reaction of the solid component on the gas, ensuring a fully coupled treatment of solid-gas interactions.
@@
In this letter, we present a comparative analysis of solid and gas torque predictions derived from simulations in canonical metallicity disks and numerical simulations in high-metallicity disk. Specifically, we contrast models assuming canonical metallicity and solid back-reaction with simulations of high-metallicity disks, in order to evaluate how metallicity and solid back-reaction interplay to influence planetary migration.
@@ 2 Hydrodynamic model
We performed two-dimensional, global hydrodynamic simulations of planet-disk interactions using the GFARGO2 code (Regály and Vorobyov, 2017; Regály, 2020), a GPU-enabled version of the FARGO code (Masset, 2000). The disk consisted of gaseous and solid components, and an embedded planet. The solid component of the disk was a pressureless fluid while the gas is locally isothermal. The dynamics of the gaseous and solid components, perturbed by the planet, are described by the continuity and Navier-Stokes equations. The dynamics of the solid component are determined by the drag force of gas arising due to the velocity difference between the two disk components. Emphasize that we also took into account the reaction force of solid component, namely the back-reaction of the solid dynamics on the gas. See Regály et al. (2025) for a complete presentation of the governing equations. To compute the dynamics of the solid, we followed the method of Stoyanovskaya et al. (2017, 2018), which was successfully tested in several works Pierens et al. (2019); Vorobyov et al. (2019); Regály et al. (2025).
We studied the effect of the solid-to-gas mass ratio, (metallicity of the disk), which is given by $\epsilon=0.01,\penalty 10000\ 0.03,\penalty 10000\$and $0.1$. For the canonical metallicity the solid-to-gas mass ratio was set to $\epsilon_{0}=0.01$. We investigated three different power law for the initial gas surface density: $p=0.5,\penalty 10000\ 1,\penalty 10000\$and $1.5$. For simplicity the solid component has a fixed Stokes number ($0.01\leq\mathrm{St}<10$) throughout the disk.
For this investigation, the planet’s mass was set to $1\penalty 10000\ M_{\oplus}$ and kept on a fixed circular orbit at $R=1$. Three models of the strengths of planetary accretion were created, assuming zero, 10 and 100 percent efficiency. Zero efficiency essentially means that there is no accretion. For moderate ($10\%$) and strong ($100\%$) accretion, the planetary Hill sphere would be emptied within 10 and 1 orbits, respectively, see details in Regály et al. (2025).
The initial conditions of the surface densities and velocities of gaseous and solid component were identical to that of applied in Regály et al. (2025). The size and geometry of the numerical grid was also identical to our previous investigation. The simulations were run for 200 planetary orbits. Emphasize that by that time both torques arising from the gas and solid components of the disk are saturated to constant values.
@@
It is natural to assume that the contribution of the solid component to the total torque scales linearly with the solid mass fraction of the disk, i.e., the metallicity. When the back-reaction of solids on the gas is neglected, this scaling is applicable, as stated in Benítez-Llambay and Pessah (2018); Regály (2020) and adopted by Guilera et al. (2023). However, once solid-gas back-reaction is included, the gas torque must be modified accordingly, with the magnitude of the correction depending on the amount of solid material. A minimal closure is to assume that the back-reaction-induced modification to the gas dynamics, and hence to the gas torque, is also linear in metallicity. Thus, to predict the solid torques for higher metallicity disk, the solid torque measured in canonical disk is simply multiplied by the metallicity increase $\Gamma^{\mathrm{BR}}_{\mathrm{d}}=\Gamma^{\mathrm{BR,\epsilon_{0}}}_{\mathrm{d}}(\epsilon/\epsilon_{0})$. However, for the predicted gas torque first we take into account of the effect of back-reaction and then amplify it by the metallicity increase $\Gamma_{\mathrm{g}}^{\mathrm{BR}}=\Gamma_{\mathrm{g}}^{\mathrm{NBR,\epsilon_{0}}}+\left(\Gamma_{\mathrm{g}}^{\mathrm{BR,\epsilon_{0}}}-\Gamma_{\mathrm{g}}^{\mathrm{NBR,\epsilon_{0}}}\right)(\epsilon/\epsilon_{0}).$ The method used to calculate the torque in the numerical simulation was the same as the one described in Regály et al. (2025).
@@ 3 Results
Figure 1: Normalized torques felt by an $1\penalty 10000\ M_{\oplus}$ planet with metallicities of $\epsilon=0.01,\penalty 10000\ 0.03,$ and $0.1$. Three accretion efficiencies were modeled: $\eta=0,\penalty 10000\ 0.1,$ and 1. The columns from left to right show three sets of models assuming different steepness of the initial disk density profile $p=0.5,1.0,$ and 1.5. Symbols represent the different Stokes numbers, $\mathrm{St}=0.01,\penalty 10000\ 0.1,\penalty 10000\ 1,\penalty 10000\ 2,\penalty 10000\ 3,\penalty 10000\ 4,\penalty 10000\ 5,$ and 10. For the elevated metallicities two models are shown: predicted (left) and numerical simulations (right). The shaded regions of positive (red), weakened negative (blue), and strengthened negative (green) torques darken with the metallicity. The normalization factor is the absolute value of the appropriate non back-reacting gas torque. Panels A1, A2, A3: solid torques in the BR models normalized by the absolute value of the solid torque NBR models having $\epsilon=0.01$. Panels B1, B2, B3: gas torques in the BR models normalized by the absolute value of the gas torque in the NBR models having $\epsilon=0.01$. Panels C1, C2, C3: total torques in the BR models normalized by the absolute value of the gas torque in the NBR models having $\epsilon=0.01$.
@@
Panels A1, A2, and A3 of Fig. 1 show the normalized torque of the solid component, $\Gamma_{\mathrm{d}}^{\mathrm{BR}}/|\Gamma_{\mathrm{g}}^{\mathrm{NBR}}|$. The subpanels, divided by the vertical black lines within each panel, represent the canonical, $\epsilon=0.03$, and $\epsilon=0.1$ metallicity cases, respectively. First, let’s discuss the canonical metallicity model, $\epsilon=0.01$, shown with the lightest shade subpanels A1a, A2a, and A3a). Note that there are no predicted torques shown for canonical model. The torque of solid component usually positive for large Stokes numbers ($\mathrm{St}\geq 2$). As the initial density profile becomes steeper, the solid torques are reduced in magnitude. Generally, the solid torque shows a weak dependence on the accretion efficiency for the canonical metallicity.
@@
Panels B1, B2, and B3 of Fig. 1 show the normalized gas torque $\Gamma_{\mathrm{g}}^{\mathrm{BR}}/|\Gamma_{\mathrm{g}}^{\mathrm{NBR}}|$. As can be seen, it is usually negative, however, exhibits weakened magnitudes for the canonical metallicity (subpanels B1a, B2a, and B3a). The degree of weakening decreases with an increasing steepness of the gas density profile. As demonstrated previously in Regály et al. (2025), the effect of solid back-reaction plays a significant role in shaping the gas torque magnitudes.
@@
Panels C1, C2, and C3 of Fig. 1 show the normalized total torque ($\Gamma_{\mathrm{g}}^{\mathrm{BR}}+\Gamma_{\mathrm{d}}^{\mathrm{BR}})/|\Gamma_{\mathrm{g}}^{\mathrm{NBR}}|$. In the canonical case (subpanels C1a, C2a, and C3a), the total torque can be positive for $p=0.5$ and large Stokes numbers, $2\leq\mathrm{St}\leq 5$ or reduced magnitude negative (compared to the non-back-reacting models) for $p=1.0$ and $p=1.5$. Note that no positive total torque was observed for $p=1.5$ models (subpanel C3a).
@@
Now, let’s compare the torque predictions and the results of the numerical simulations for three times the canonical metallicity, i.e. for a value of the parameter $\epsilon=0.03$. The predicted torques of the solid components practically match those computed from the simulations (subpanels A1b, A2b, and A3b). However, the predicted gas torques differ quantitatively from the predictions (subpanels B1b, B2b, and B3b). Note that when making the prediction, we simply assumed that the effect of back-reaction was three times that observed in the canonical metallicity model. The largest discrepancy between the predicted gas torque and that computed from numerical simulations is about 50% for $1\leq\mathrm{St}\leq 5$. For the coupled species ($\mathrm{St}\leq 0.1$), the discrepancy is modest. Consequently, the total torques derived from numerical simulations can be accurately predicted for $\mathrm{St}\leq 0.1$ (see, e.g., $\mathrm{St}=0.01$, C1b). However, for $1\leq\mathrm{St}\leq 5$, the total torque magnitudes can be overpredicted by up to 100%, as can be observed for $2\leq\mathrm{St}\leq 3$ (e.g., subpanel C1b). It should be emphasized that the above findings are not affected by accretion efficiency.
@@
Now, let’s focus on models that use ten times the canonical metallicity, i.e. with $\epsilon=0.1$. Unlike the models with a metallicity of 0.03, in this case the predicted and numerical torques of the solid component show significant quantitative differences for $1\leq\mathrm{St}\leq 3$. The torques arising from these species are weaker or stronger in the numerical models than in the predictions, when they are negative or positive, respectively. A much larger discrepancy is found with regard to the gas torque (subpanels B1c, B2c, and B3c). For $\mathrm{St}=0.01$ and $\mathrm{St}=1$, the sign of the torque can even mismatch. For instance, in the $\mathrm{St}=0.01$ and $p=0.5$ or $p=1.0$ models (see $\mathrm{St}=0.01$ in subpanels B1c and B2c), the torque predictions are positive for accreting planets with high accretion efficiency, but the numerical simulations give strong negative torques. Another example is the $\mathrm{St}=1$ and $p=0.5$ model (subpanel B1c, $\mathrm{St}=1$), where the predictions give strong negative torques, while the numerical simulations result in significant positive torques, independent of the accretion efficiency. Comparing the total torques (see C1c, C2c, and C3c subpanels of Fig. 1), we find that the predictions and numerical simulations agree well for $\mathrm{St}>2$, while significant discrepancies are found for $\mathrm{St}\leq 2$. Moreover, the total torques arising in $\mathrm{St}=0.01$ and $p=0.5$ and $p=1.0$ or $\mathrm{St}=2$ and $p=1.0$ models, can differ in sign due to the combined effect of changes in the solid and gas torques (see, e.g., $\mathrm{St}=2$ in subpanel C2c).
@@
To understand the origin of these discrepancies, we analyze the gas surface density perturbations induced by the planet in both canonical and high-metallicity disk simulations (see Section 4 for details). Our analysis indicates that the qualitative disagreement between the predicted and simulated torques arises from a strongly asymmetric perturbation in the co-orbital region, with respect to the planet’s orbital motion.

### v2

@@ Abstract
Context. The migration of low‑mass planets ($M_{\mathrm{p}}\leq 10\penalty 10000\ M_{\oplus}$) is tightly controlled by the torques exerted by both the gas and solids in their natal disks. While canonical models assume a solar solid‑to‑gas mass ratio ($\epsilon\simeq 0.01$) and neglect the back-reaction of the solid component of the disk, recent work suggests that enhanced metallicity can radically alter these torques.
@@
Methods. We performed global, two‑dimensional hydrodynamic simulations that (i) treat solid material as a pressureless fluid fully coupled to the gas through drag and (ii) include the reciprocal back‑reaction force. The low-mass planet was maintained on a fixed circular orbit, and thus we computed static torques. The Stokes number was varied from 0.01 to 10, and three surface‑density slopes ($p=0.5,\penalty 10000\ 1.0,\mathrm{and}\penalty 10000\ 1.5$) and three accretion efficiencies ($\eta=0\%,\penalty 10000\ 10\%,\mathrm{and}\penalty 10000\ 100\%$) were explored. Predicted torques, obtained by rescaling canonical $\epsilon=0.01$ results, were compared with direct simulations.
Results. Solid torques scale nearly linearly with $\epsilon$, but gas torques deviate by $50-100\%$ and can even reverse sign for $\mathrm{St}\leq 1$ in $\epsilon=0.1$ disks. These discrepancies arise from strong, feedback‑driven, asymmetric gas perturbations in the co‑orbital region, which are amplified by rapid planetary accretion. Accurate total torques are recovered only for $\mathrm{St}\geq 3$, independent of $\epsilon$ or $\eta$; for $\mathrm{St}\leq 2$ the linear prescription systematically overestimates the magnitude, sometimes predicting the wrong sign.
@@ 1 Introduction
It is well established that planets undergo changes in their semimajor axes due to gravitational interactions with their natal protoplanetary disk. In particular, low-mass planets (with masses $\lesssim 10\penalty 10000\ M_{\oplus}$) are generally subject to inward migration in locally isothermal disks as a result of disk torques (Tanaka et al., 2002). Consequently, planets that remain embedded in the disk during its lifetime are at risk of spiraling into their host star.
@@
Several mechanisms that potentially counteract this potentially catastrophic inward migration have been proposed, including the concept of planetary migration traps (Lyra et al., 2008; Sándor et al., 2011; Guilera and Sándor, 2017; Pierens and Raymond, 2024) and localized temperature perturbations within the disk (Lyra et al., 2010). For comprehensive reviews of the physical processes governing planetary migration, see, for example, Kley and Nelson (2012), Raymond et al. (2020), and Paardekooper et al. (2023).
@@
In recent years, numerical simulations of planet–disk interactions have introduced a new paradigm, emphasizing the pivotal role of the disk’s solid component in altering – and in some cases reversing – the direction of planetary migration. It has been demonstrated that a pronounced asymmetry in the spatial distribution of solids near the planet can generate an additional torque, which may reach magnitudes comparable to that of the gas torque (Benítez-Llambay and Pessah, 2018). Moreover, planetary accretion can further enhance this solid-induced torque, thereby amplifying its impact on the planet’s orbital evolution (Regály, 2020; Chrenko et al., 2024; Pierens, 2023). When these findings are applied within a global disk model that incorporates viscous accretion and X-ray photoevaporation of the gaseous component, along with dust evolution processes, it becomes evident that the torque exerted by the solid phase can significantly influence migration behavior (Guilera et al., 2023). In particular, such models reveal that planets forming near the water-ice line can undergo substantial outward migration driven by solid torques (Guilera et al., 2025).
@@
The back-reaction of solid material on the gas is often neglected in numerical simulations of planet–disk interactions. Nonetheless, a growing body of work has underscored the significance of this feedback mechanism (Gonzalez et al., 2017; Dipierro et al., 2018; Gonzalez et al., 2018; Hsieh and Lin, 2020). Our recent study has shown that back-reaction from the solid component can substantially perturb the gas distribution in the vicinity of the planet, to the extent that the gas torque becomes positive, even under locally isothermal conditions (Regály et al., 2025). In addition, a recent three-dimensional hydrodynamic simulation demonstrated that this solid back-reaction can excite buoyancy resonances in a globally isothermal disk that would otherwise lack such instabilities (Chametla et al., 2025b).
@@
In the canonical scenario, the solid-to-gas mass ratio in protoplanetary disks is approximately one percent (e.g., Williams and Cieza, 2011). It is commonly assumed that the torque exerted by the solid component scales linearly with the total solid mass in the disk (see, e.g., Benítez-Llambay and Pessah 2018). This assumption may break down when the back-reaction from the solid component becomes sufficiently strong to perturb the gas dynamics. Such perturbations can significantly affect the evolution of the solid component itself, resulting in a nonlinear and coupled torque response.
@@
Hsieh and Lin (2020) show that higher-metallicity disks can produce solid torques strong enough to alter the migration direction of low-mass planets or to induce stochastic migration behavior. The stochastic nature of migration is associated with the excitation of small-scale vortices produced in inviscid simulations. Their work also accounted for the dynamical back-reaction of the solid component on the gas, ensuring a fully coupled treatment of solid–gas interactions.
@@
We present a comparative analysis of solid and gas torque predictions derived from simulations in canonical-metallicity disks and numerical simulations in high-metallicity disks. Specifically, we contrast models that assume a canonical metallicity and solid back-reaction with simulations of high-metallicity disks in order to evaluate how metallicity and solid back-reaction interplay to influence planetary migration.
@@ 2 Hydrodynamic model
We performed two-dimensional, global hydrodynamic simulations of planet-disk interactions using the GFARGO2 code (Regály and Vorobyov, 2017; Regály, 2020), a GPU-enabled version of the FARGO code (Masset, 2000). The disk consisted of gaseous and solid components, and an embedded planet. The solid component of the disk was a pressureless fluid while the gas is locally isothermal. The dynamics of the gaseous and solid components, perturbed by the planet, are described by the continuity and Navier-Stokes equations. The dynamics of the solid component are determined by the drag force of gas arising due to the velocity difference between the two disk components. Emphasize that we also took into account the reaction force of solid component, namely the back-reaction of the solid dynamics on the gas. See Regály et al. (2025) for a complete presentation of the governing equations. To compute the dynamics of the solid, we followed the method of Stoyanovskaya et al. (2017, 2018), which was successfully tested in several works (Pierens et al., 2019; Vorobyov et al., 2019; Regály et al., 2025).
We studied the effect of the solid-to-gas mass ratio (metallicity of the disk), which is given by $\epsilon=0.01,\penalty 10000\ 0.03,\penalty 10000\$and $0.1$. For the canonical metallicity, the solid-to-gas mass ratio was set to $\epsilon_{0}=0.01$. We investigated three different power law for the initial gas surface density: $p=0.5,\penalty 10000\ 1,\penalty 10000\$and $1.5$. For simplicity the solid component has a fixed Stokes number ($0.01\leq\mathrm{St}<10$) throughout the disk.
For this investigation, the planet’s mass was set to $1\penalty 10000\ M_{\oplus}$ and kept on a fixed circular orbit at $R=1$. Three models of the strengths of planetary accretion were created, assuming zero, 10 and 100 percent efficiency. Zero efficiency essentially means that there is no accretion. For moderate ($10\%$) and strong ($100\%$) accretion, the planetary Hill sphere would be emptied within 10 and 1 orbits, respectively; see details in Regály et al. (2025).
The initial conditions of the surface densities and velocities of gaseous and solid component were identical to that of applied in Regály et al. (2025). The size and geometry of the numerical grid was also identical to our previous investigation. The simulations were run for 200 planetary orbits. We emphasize that, by that time, the torques arising from both the gas and solid components of the disk are saturated to constant values.
@@
It is natural to assume that the contribution of the solid component to the total torque scales linearly with the solid mass fraction of the disk, i.e., the metallicity. When the back-reaction of solids on the gas is neglected, this scaling is applicable, as stated in Benítez-Llambay and Pessah (2018); Regály (2020) and adopted by Guilera et al. (2023). However, once solid-gas back-reaction is included, the gas torque must be modified accordingly, with the magnitude of the correction depending on the amount of solid material. A minimal closure is to assume that the back-reaction-induced modification to the gas dynamics, and hence to the gas torque, is also linear in metallicity. Thus, to predict the solid torques for higher metallicity disk, the solid torque measured in canonical disk is simply multiplied by the metallicity increase $\Gamma^{\mathrm{BR}}_{\mathrm{d}}=\Gamma^{\mathrm{BR,\epsilon_{0}}}_{\mathrm{d}}(\epsilon/\epsilon_{0})$. However, for the predicted gas torque, first we took the effect of the back-reaction into account and then amplified it by the metallicity increase $\Gamma_{\mathrm{g}}^{\mathrm{BR}}=\Gamma_{\mathrm{g}}^{\mathrm{NBR,\epsilon_{0}}}+\left(\Gamma_{\mathrm{g}}^{\mathrm{BR,\epsilon_{0}}}-\Gamma_{\mathrm{g}}^{\mathrm{NBR,\epsilon_{0}}}\right)(\epsilon/\epsilon_{0}).$ The method used to calculate the torque in the numerical simulation was the same as the one described in Regály et al. (2025).
@@ 3 Results
[…]
@@
Panels A1, A2, and A3 of Fig. 1 show the normalized torque of the solid component, $\Gamma_{\mathrm{d}}^{\mathrm{BR}}/|\Gamma_{\mathrm{g}}^{\mathrm{NBR}}|$. The subpanels, divided by the vertical black lines within each panel, represent the canonical, $\epsilon=0.03$, and $\epsilon=0.1$ metallicity cases, respectively. First, let us discuss the canonical-metallicity model, $\epsilon=0.01$, shown with the lightest shade subpanels A1a, A2a, and A3a). Note that there are no predicted torques shown for canonical model. The torque of solid component usually positive for large Stokes numbers ($\mathrm{St}\geq 2$). As the initial density profile becomes steeper, the solid torques are reduced in magnitude. Generally, the solid torque shows a weak dependence on the accretion efficiency for the canonical metallicity.
@@
Panels B1, B2, and B3 of Fig. 1 show the normalized gas torque, $\Gamma_{\mathrm{g}}^{\mathrm{BR}}/|\Gamma_{\mathrm{g}}^{\mathrm{NBR}}|$. As can be seen, it is usually negative; however, it exhibits weakened magnitudes for the canonical metallicity (subpanels B1a, B2a, and B3a). The degree of weakening decreases with an increasing steepness of the gas density profile. As demonstrated previously in Regály et al. (2025), the effect of solid back-reaction plays a significant role in shaping the gas torque magnitudes.
@@
Panels C1, C2, and C3 of Fig. 1 show the normalized total torque, ($\Gamma_{\mathrm{g}}^{\mathrm{BR}}+\Gamma_{\mathrm{d}}^{\mathrm{BR}})/|\Gamma_{\mathrm{g}}^{\mathrm{NBR}}|$. In the canonical case (subpanels C1a, C2a, and C3a), the total torque can be positive for $p=0.5$ and large Stokes numbers, $2\leq\mathrm{St}\leq 5$ or reduced magnitude negative (compared to the non-back-reacting models) for $p=1.0$ and $p=1.5$. Note that no positive total torque was observed for $p=1.5$ models (subpanel C3a).
@@
Now, let us compare the torque predictions and the results of the numerical simulations for three times the canonical metallicity, i.e., for a value of the parameter $\epsilon=0.03$. The predicted torques of the solid components practically match those computed from the simulations (subpanels A1b, A2b, and A3b). However, the predicted gas torques differ quantitatively from the predictions (subpanels B1b, B2b, and B3b). Note that when making the prediction, we simply assumed that the effect of back-reaction was three times that observed in the canonical-metallicity model. The largest discrepancy between the predicted gas torque and that computed from numerical simulations is about 50% for $1\leq\mathrm{St}\leq 5$. For the coupled species ($\mathrm{St}\leq 0.1$), the discrepancy is modest. Consequently, the total torques derived from numerical simulations can be accurately predicted for $\mathrm{St}\leq 0.1$ (see, e.g., $\mathrm{St}=0.01$, C1b). However, for $1\leq\mathrm{St}\leq 5$, the total torque magnitudes can be over-predicted by up to 100%, as can be observed for $2\leq\mathrm{St}\leq 3$ (e.g., subpanel C1b). It should be emphasized that the above findings are not affected by accretion efficiency.
@@
Now, let us focus on models that use ten times the canonical metallicity, i.e., with $\epsilon=0.1$. Unlike the models with a metallicity of 0.03, in this case the predicted and numerical torques of the solid component show significant quantitative differences for $1\leq\mathrm{St}\leq 3$. The torques arising from these species are weaker or stronger in the numerical models than in the predictions, when they are negative or positive, respectively. A much larger discrepancy is found with regard to the gas torque (subpanels B1c, B2c, and B3c). For $\mathrm{St}=0.01$ and $\mathrm{St}=1$, the sign of the torque can even mismatch. For instance, in the $\mathrm{St}=0.01$ and $p=0.5$ or $p=1.0$ models (see $\mathrm{St}=0.01$ in subpanels B1c and B2c), the torque predictions are positive for accreting planets with high accretion efficiency, but the numerical simulations give strong negative torques. Another example is the $\mathrm{St}=1$ and $p=0.5$ model (subpanel B1c, $\mathrm{St}=1$), where the predictions give strong negative torques, while the numerical simulations result in significant positive torques, independent of the accretion efficiency. Comparing the total torques (see the C1c, C2c, and C3c subpanels of Fig. 1), we find that the predictions and numerical simulations agree well for $\mathrm{St}>2$, while significant discrepancies are found for $\mathrm{St}\leq 2$. Moreover, the total torques arising in $\mathrm{St}=0.01$ and $p=0.5$ and $p=1.0$ or $\mathrm{St}=2$ and $p=1.0$ models, can differ in sign due to the combined effect of changes in the solid and gas torques (see, e.g., $\mathrm{St}=2$ in subpanel C2c).
@@
To understand the origin of these discrepancies, we analyzed the gas surface density perturbations induced by the planet in both canonical and high-metallicity disk simulations (see Sect. 4 for details). Our analysis indicates that the qualitative disagreement between the predicted and simulated torques arises from a strongly asymmetric perturbation in the co-orbital region, with respect to the planet’s orbital motion.